how do I find the area of a running track shaped like a rectangle with a semicircle on each end. The distance around the track is 5280 ft. The straightaway is twice the width of the field I need to find the area enclosed by the track to the nearest sq foot. What formula would I use?
Answer by checkley77(12844)

FIND THE WIDTH OF THE FIELD (X), THE LENGTH OF THE STRAIGHTAWAY (2X) ANS THE
CIRCUMFERANCE OF THE SEMICIRCLES (2PIR) AND THE AREA OF THE CIRCLE (PIR^2).
-----------------------
STRAIGHT AWAY=2*2X=4X
END SEMICIRCLES=2PIX=2*3.14X=6.28X
4X+6.28X=5,280
10.28X=5,280
X=5,280/10.28
X=513.62 IS THE WIDTH OF THE FIELD.
2*513.62=1,027.24 LENGTH OF THE STRAIGHTAWAY.
2*3.14*513.62=6.28*513.62=3,225.53 DISTANCE AROUND THE 2 SEMI CIRCLES.
PROOF:
2*1,027.24+3,225.53=2054.48+3,225.53=5,280 DISTANCE OF THE TRACK.
3.14*513.62^2=3.14*263,805.5 AREA OF THE END SEMI CIRCLES.
496.8*939.6=466,793.28 AREA OF THE RECTANGLE.
466,793.28+233,396.64=700,189.92 FT^2 ANS. FOR THE AREA OF THE FIELD.